2. Build a voltaic cell with one beaker containing a solution of nickel (II) nitrate - Ni(NO3)2

- with a nickel electrode, and a second beaker containing a solution of zinc chloride - 

ZnCl2 - with a zinc electrode.

a.  The half reactions

b.  Flow of electrons

c.  Short-form notation

d.  Electric potential (voltage)


1
Expert's answer
2021-02-03T07:15:29-0500

a). Zn + 2Cl- = ZnCl2 + 2e-

Ni(NO3)2 + 2e- = Ni2+ + 2(NO3)2

b). Zn = Zn2+ + 2e-

Ni = Ni2+ + 2e-

Zn + Ni2+ "\to" Zn2+ + Ni

c). Zn(s) ZnCl2(s) Cl-(aq) Ni(NO3)2(s) Ni(s)

d). "\phi" (Zn) = -0.762B

"\phi" (Ni) = -0.257B

0.762+0.257 = 1.019B